4. Calcium carbonate (CaCO3) which is found in limestone and marble reacts readily with hydrochloric acid (HCl) to produce calcium chloride (CaCl2), carbon dioxide, and water.


a. Write the balanced chemical reaction:


b. What mass of hydrochloric acid will be needed to produce 325 g of calcium chloride?


c. What mass of carbon dioxide could be produced if 750 g of calcium carbonate reacts with excess hydrochloric acid?



1
Expert's answer
2021-05-27T07:36:47-0400

Solution:

a)

Balanced chemical equation:

CaCO3(s) + 2HCl(aq) → CaCl2(aq) + CO2(g) + H2O(l)


b)

The molar mass of HCl is 36.458 g/mol.

The molar mass of CaCl2 is 110.98 g/mol.

According to the equation above: n(HCl)/2 = n(CaCl2)

Therefore,

(325 g CaCl2) × (1 mol CaCl2 / 110.98 g CaCl2) × (2 mol HCl / 1 mol CaCl2) × (36.458 g HCl / 1 mol HCl) = 213.53 g HCl


c)

The molar mass of CaCO3 is 100.087 g/mol.

The molar mass of CO2 is 44.01 g/mol.

According to the equation above: n(CaCO3) = n(CO2)

Therefore,

(750 g CaCO3) × (1 mol CaCO3 / 100.087 g CaCO3) × (1 mol CO2 / 1 mol CaCO3) × (44.01 g CO2 / 1 mol CO2) = 329.79 g CO2


Answers:

a) CaCO3(s) + 2HCl(aq) → CaCl2(aq) + CO2(g) + H2O(l)

b) 213.53 g HCl

c) 329.79 g CO2